OpenStudy (anonymous):

you flip three coins. what is the probability that you get at least two tails, given that you get at least one tail?

OpenStudy (anonymous):

given you have at least one tail cuts down the sample space, and gets rid of the even that you get (h, h, h) so you have 7 elements in the sample space, rather than the usual 8

OpenStudy (anonymous):

total sample space is now (h, h , t) (h, t, h) (t, h, t) (h, t, t) (t, h, t) (t, t, h) (t, t, t) of those you can count how many have at least two tails

OpenStudy (anonymous):

sorry i meant "gets rid of the EVENT that you get 3 heads"

OpenStudy (anonymous):

i still don't get it...

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

the line that reads " given that you get at least one tail?" means you know for sure you have tossed at least one tail in the three tosses

OpenStudy (anonymous):

i got that

OpenStudy (anonymous):

when you flip 3 coins, since there are two possible outcomes for each coin (H or T) you know by the counting principle that there are \(2\times 2\times 2=2^3=8\) possible outcomes since 8 is not such a big number, we can list all of them

OpenStudy (anonymous):

(h, h, h) (h, h , t) (h, t, h) (t, h, t) (h, t, t) (t, h, t) (t, t, h) (t, t, t) there are the 8 possibilities for tossing 3 coins

OpenStudy (anonymous):

of those 8, since you are told you have at least one tail, we can get rid of the first one, since that one has not tails

OpenStudy (anonymous):

*no tails

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

that leaves 7 possible outcomes, not 8

OpenStudy (anonymous):

now you want to count, out of those 7, how many have "at least 2 tails"

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

i count exactly 4, namely (h, t, t) (t, h, t) (h, t, t) and (t, t, t)

OpenStudy (anonymous):

oh yea ok 4

OpenStudy (anonymous):

those have "at least two tails"

OpenStudy (anonymous):

so the answer is four out of seven, i.e. \(\frac{4}{7}\)

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

yw